# DarkAI: Reconstructing the density, velocity and tidal field of dark matter from DESI-like bright galaxy sample

### v1

@@ 3 METHODOLOGY
3.1 Reconstruction via Maximum-Likelihood Estimation
Reconstructing the dark matter density and velocity fields from the galaxy density field constitutes a field-to-field mapping problem, which can be framed as a maximum likelihood estimation (MLE) task (e.g. Mao et al., 2021). The objective is to infer the underlying dark matter fields given the observed galaxy density field. Assuming a probabilistic relationship between the two, the likelihood function is given by,
@@ 3.1 Reconstruction via Maximum-Likelihood Estimation
The estimation of $\boldsymbol{\theta}$ can be expressed in the following log-likelihood form,
[…]
@@
However, solving this formulation analytically is impractical due to the model’s complexity. Traditional optimization methods face challenges in estimating $\boldsymbol{\theta}$ due to the high dimensionality of the parameter space. Alternatively, neural networks provide an efficient approach by maximizing likelihood through supervised learning, effectively encoding $\boldsymbol{\theta}$ in their weights and biases.
Figure 4: Three-dimensional (3D) UNet architecture. Each cube represents a 3D space-with-feature map, with the grid size and the number of feature channels indicated. The input and output cubes, each containing $512^{3}$ voxels and spanning $1564\,\>h^{-1}{\rm{Mpc}}$, correspond to the redshift-space galaxy density field and the real-space dark matter density or velocity field, respectively. The encoder comprises 8 convolutional layers that progressively downsample the input cube to a bottleneck representation, while the decoder uses 8 transposed convolutional layers to upsample the bottleneck output to the target cube. The channel configuration (1+3) corresponds to the density field, $\delta_{\mathrm{m}}$, and the velocity components $(v_{x},v_{y},v_{z})$, enabling simultaneous reconstruction of these quantities as an optional feature. Computational steps are represented by distinct symbols, as illustrated in the lower-right corner. All convolutional layers utilize $4\times 4\times 4$ filters with a stride of $2$.
@@ 3.2 UNet Architecture
We base our network architecture on the U-Net style (Ronneberger et al., 2015), which is particularly well-suited for reconstructing complex fields due to its hierarchical convolutional layers and skip connections. These design features allow the model to capture multi-scale spatial information effectively, which is critical for learning the mapping between the redshift-space galaxy density field and the target fields.
@@
  • Encoder Path: The encoder extracts features from the input field, $\delta_{\mathrm{g}}$, while progressively reducing spatial resolution. Each block consists of a sequence: convolution, batch normalization, and LeakyReLU activation. Batch normalization is omitted in the first and last layers of the encoder, where the last layer uses ReLU activation instead of LeakyReLU.
  • Decoder Path: The decoder reconstructs the target field incrementally by increasing spatial resolution. Each block consists of a sequence: transposed convolution, batch normalization, dropout (with a rate of 50%), and ReLU activation. Batch normalization is not applied in the first layer of the decoder, and the output layer employs a linear activation function.
  • Skip Connections: Skip connections link encoder layers to their corresponding decoder layers, enabling the preservation of spatial details during downsampling and upsampling.
  • Output Layer: For density field reconstruction, the output layer generates a single-channel output corresponding to $\delta_{\mathrm{m}}$. For velocity field reconstruction, the output layer produces three channels, representing the velocity components $(v_{x},v_{y},v_{z})$.
In this architecture, the fields $\delta_{\mathrm{m}}$ and $\boldsymbol{v}_{\mathrm{m}}$ are reconstructed separately due to memory limitations, while they could be reconstructed simultaneously by designing a multichannel output for the network if hardware permits.
Figure 5: (Upper left) Loss function for training $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ using 3000 epochs. The dark solid and blue dashed lines represent the loss functions from the training and validation samples, respectively. (Upper right) Comparison of the density distributions for the validation samples. The black solid line represent the result of the target validation samples. Different colored lines represent the predicted fields at epochs 200, 1000, 2000, and 3000, repectively. (Lower left) Grid-to-grid density relation between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$. The contour in each case represents threshold encompassing 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. Different colored curves correspond to different epochs, as indicated. (Lower right) The evolution of the factor $f_{sc}$, defined as the product of the scatter variance and one minus the correlation scope between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as a function of epoch.
@@ 3.3 Training Process
As described in Section 2, we generate paired data samples $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ and $(\delta_{\mathrm{g}},\boldsymbol{v}_{\mathrm{m}})$ to train the networks. These samples represent the galaxy density field $\delta_{\mathrm{g}}$ paired with the true dark matter density field $\delta_{\mathrm{m}}$ or the velocity field $\boldsymbol{v}_{\mathrm{m}}$. The objective is to learn the mapping between the galaxy and dark matter fields.
The training process then minimizes the MSE loss function, which corresponds to the log-likelihood function in Equation (5). For practical implementation, the loss function is expressed as,
@@
where $\phi^{\mathrm{pred}}_{\mathrm{m}}$ is the field being reconstructed and $\phi^{\mathrm{true}}_{\mathrm{m}}$ denote the true target field. Specifically, $\phi_{\mathrm{m}}\equiv\delta_{\mathrm{m}}$ for density reconstruction and $\phi_{\mathrm{m}}\equiv\boldsymbol{v}_{\mathrm{m}}$ for velocity reconstruction.
To improve the training stability and optimize the network’s ability to capture variations across a wide range of values, we rescale the input fields before training. The overdensity field is rescaled using a logarithmic transformation,
@@
where the parameter $a=0.02$ ensures numerical stability by preventing singularities for low-density regions ($\delta\approx-1$). Moreover, the $a$ value is optimized in our analysis to balance the reconstruction, enhancing accuracy in high-density regions while preserving reliable performance in low-density areas. The velocity field is rescaled linearly as,
@@
where the parameter $b=3000\,\mathrm{km}\,\mathrm{s}^{-1}$ normalizes the velocity values to a comparable scale, avoiding excessively large gradients during training. These rescaling transformations enhance the network’s convergence and improve its performance by standardizing the input dynamic range.
@@
The network is trained from scratch using the Adam optimizer (Loshchilov & Hutter, 2017), which incorporates the adaptive learning rate. Each epoch processes 5 training samples, followed by 2 validation samples used exclusively for model evaluation during training.
@@
In training neural networks, selecting an optimal stopping point is critical to balancing generalization and overfitting, particularly for fields with wide-range density distributions. As shown in the upper-left panel of Figure 5, the validation loss converges around epoch 200, indicating an appropriate point to terminate training. However, relying solely on loss metrics may overlook model performance across varying density regions, especially in high-density areas, which represent a small fraction of the total volume.
@@
To address this limitation, we compare the density distributions of the predicted and true fields (upper-right panel of Figure 5). From epoch 200 to 2000, the disparities in high-density regions with $(1+\delta)>6$ progressively diminish, indicating significant improvements in recovering high-density structures even after loss convergence. However, as training progresses, predictions exhibit systematic underestimation at epoch 3000, revealing overfitting effects.
@@
However, the grid-to-grid relationship (lower-left panel of Figure 5) support the opposite trend. From epoch 200 to 2000, the scatter increases while maintaining minimal bias, indicating improved density distribution recovery at the cost of reduced correlation with the true field. This trade-off highlights the need for a quantitative criterion to identify the optimal training stage.
@@
We define the factor $f_{sc}$ as the product of the perpendicular variance, represented by the smaller eigenvalue of the covariance matrix of $(\vec{\delta}_{\mathrm{rec}},\vec{\delta}_{\mathrm{true}})$, and the absolute value of one minus the correlation slope, which is given by the slope of the eigenvector corresponding to the larger variance. As illustrated in the lower-right panel of Figure 5, this factor reaches its minimum value at approximately epoch 1300 and increases significantly after epoch 2000. Consequently, halting the training process around epoch 1300 achieves a balance between high-density recovery and correlation retention, ensuring the reconstructed field accurately represents cosmological structures.
@@
These results underscore the necessity of carefully assessing the training stages to ensure the reliability of the model for cosmological inference. The methodology introduced here could be generalized to other datasets, offering a pathway to robust machine learning applications in astrophysical studies.
@@
Figure 6: Comparison of the projected density-field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. Different panels correspond to the redshift-space galaxy distribution (input), the true dark matter density field $\delta_{\mathrm{true}}$ (target), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (output), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as labeled at the top of each panel.
@@
Figure 7: (Left panel) Grid-to-grid density relationship between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$, represented by the blue solid line. For comparison, the result for the galaxy density field is also plotted with the red dashed line. The three contours in each case represent thresholds encompassing 67%, 95%, and 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. (Right panel) Density distributions for galaxies (red dashed), the reconstructed field (blue solid), and the true field (black solid). Error bars indicate the ±$\pm 1\sigma$ variance among the 5 test samples.
@@
Figure 8: Same as Figure 7 but for comparing the densities in low-redshift (orange) and high-redshift region (blue).
@@
Figure 9: Comparison of $P(k)$ (upper panel) and $P(k)$ ratios (lower panel). The black solid line and blue dashed line represent the auto-power spectrum of the true and reconstructed density fields respectively, while the green dot-dashed line represents their cross-power spectra. For reference, the red dotted line shows the auto-power spectra of the redshift-space galaxy field. The bottom panel shows the ratio between the reconstructed and true power spectra, with error bars indicating the ±$\pm\sigma$ variance across the five test samples. The black solid error bars are computed using Eq. (9).
@@
Figure 10: The ratios $R_{\ell}(k)$ for the quadrupole (blue solid) and hexadecapole (orange dashed) power spectrum.
@@ 4 Results
In this section, we present the reconstructed density and velocity fields derived from the mock DESI survey by applying the resulting UNet model to 5 test samples. We also assess the reconstructed density field by validating its tidal field.
4.1 Reconstructing density field
For visual comparison, Figure 6 shows the projected density-field distribution in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. Different panels correspond to the redshift-space galaxy distribution (upper left), the true dark matter density field $\delta_{\mathrm{true}}$ (upper right), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (lower left), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$ (lower right). Compared with the true field, the reconstructed density field reveals recognizable and consistent large-scale structures, including clusters, filaments, and voids, indicating a successful reconstruction overall. Additionally, the results demonstrate resilience to boundary effects, as structures near the survey edges are effectively recovered. However, the difference map highlights discrepancies in high-density regions, particularly around clusters, where the reconstructed field deviates from the target.
@@ 4.1 Reconstructing density field
The right panel of Figure 7 shows the density distributions for galaxies (red dashed), the reconstructed field (blue solid), and the true field (black solid). The error bars indicate the ±$\pm 1\sigma$ variance among the 5 test samples. After reconstruction, $\delta_{\mathrm{rec}}$ closely matches $\delta_{\mathrm{true}}$ in the regions where $71+\delta<$ 7, but a systematic underestimation occurs for $71+\delta>$ 7. This discrepancy likely arises from the imbalance in the number of grid cells between low- and high-density regions. As discussed in Section 3.3, the network initially prioritizes accurate predictions in low-density regions during early training stages, gradually improving high-density recovery at the cost of reduced correlation with the true field in later stages. The final model was selected by balancing scatter and correlation measures, which may contribute to the observed divergence. However, this issue affects only about 0.2% of the grids with the $71+\delta>$ 7 region. In comparison, the galaxy density distribution fails to trace the dark matter field accurately across both low- and high-density regimes.
@@
Additionally, Figure 8 presents the comparison between the low-redshift ($0.1\lesssim z\lesssim 0.2$) and high-redshift ($0.3\lesssim z\lesssim 0.4$) regions. In the left panel, the high-redshift region exhibits slightly larger scatter, highlighting the impact of lower galaxy density. In the right panel, the reconstructed density distribution underestimates high-density regions more significantly in the low-redshift region. This underestimation likely arises from higher galaxy clustering at low redshifts, which enhances nonlinear effects that the reconstruction method cannot fully correct, resulting in a suppression of high-density peaks. Overall, the method demonstrates robust performance, though further refinement is needed to better address nonlinear effects.
@@
To evaluate the clustering, we compare the power spectrum of the reconstructed dark matter density fields with the true fields. In this study, random samples were not used to correct for the survey’s geometry or selection effects, as the focus is on evaluating the reconstruction’s fidelity through direct comparison. This highlights the quality of the reconstruction but does not represent the true underlying clustering.
@@
The top panel of Figure 9 illustrates the power spectrum comparisons. The black solid line and blue dashed line represent the auto-power spectra of the true and reconstructed density fields, respectively. For reference, the red dotted line shows the auto-power spectrum of the redshift-space galaxy density field. The bottom panel shows the ratio between the reconstructed and true power spectra, with error bars indicating the ±$\pm\sigma$ variance across the five test samples. By considering the cosmic variance due to the limited number of Fourier modes for a finite-volume survey, we presented the theoretical error (black error bar) for the power spectrum(e.g. Song et al., 2021) by,
@@
Figure 11: Comparison between the true and reconstructed dark-matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented.
To evaluate the concordance between the reconstructed and true density fields, we analyzed the cross-correlation power spectrum. Unlike the auto-power spectrum, which assesses the overall clustering amplitude, the cross-power spectrum directly measures how well the reconstructed field reproduces the phase and spatial distribution of the true field. High cross-correlation values indicate that the reconstruction effectively captures both large- and small-scale structures, making it a robust validation metric. In Figure 9, the green dash-dotted line represents the cross-power spectrum of the true and reconstructed fields, as well as the ratios of the reconstructed cross-power spectrum to the true one. The results show that the ratios align with unity within the $1\sigma$ range on scales $k<0.1\,h\,\mathrm{Mpc}^{-1}$, indicating a strong agreement between the two fields. At smaller scales ($k>0.1\,h\,\mathrm{Mpc}^{-1}$), the ratios fall below unity. Specifically, the reductions in the cross-correlation power spectrum are approximately 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$. These reductions, though moderate, reflect challenges in accurately reconstructing high-density regions at smaller scales. Quantitatively, the average cross-correlation coefficient remains above 0.985 for $k<0.1\,h\,\mathrm{Mpc}^{-1}$, confirming the high fidelity of the reconstructed field for large-scale structures.
To evaluate the accuracy of redshift-space distortion (RSD) corrections in the reconstructed dark matter density field, we compute the ratios of the multipole moments of the power spectra. These ratios compare the reconstructed and true power spectra, normalized by the difference between the redshift-space galaxy density power spectrum and the true power spectrum. The general formula for the ratio for any multipole $\ell$ is expressed as:

### v2

@@ 3 METHODOLOGY
3.1 Reconstruction via Maximum Likelihood Estimation
Reconstructing the mass density and velocity fields from the galaxy density field constitutes a field-to-field mapping problem, which can be framed as a maximum likelihood estimation task (e.g. Mao et al., 2021). The objective is to infer the underlying dark matter fields, given the observed galaxy density field. Assuming a probabilistic relationship between the two, the likelihood function is given by
@@ 3.1 Reconstruction via Maximum Likelihood Estimation
The estimation of $\boldsymbol{\theta}$ can be expressed in the following log-likelihood form:
$\hat{\boldsymbol{\theta}}_{\mathrm{ML}}=\underset{\theta}{\mathrm{arg\,min}}\sum_{k}\left[f(\delta_{\mathrm{g},k};\boldsymbol{\theta})-\phi_{\mathrm{m},k}\right]^{2},$ (5)
@@
However, solving this formulation analytically is impractical, due to the model’s complexity. Traditional optimization methods face challenges in estimating $\boldsymbol{\theta}$ due to the high dimensionality of the parameter space. Alternatively, neural networks provide an efficient approach by maximizing likelihood through supervised learning, effectively encoding $\boldsymbol{\theta}$ in their weights and biases.
Figure 5: 3D UNet architecture. Each cube represents a 3D space-with-feature map, with the grid size and the number of feature channels indicated. The input and output cubes, each containing $512^{3}$ voxels and spanning $1564\,\>h^{-1}{\rm{Mpc}}$, correspond to the redshift-space galaxy density field and the real-space dark matter density or velocity field, respectively. The encoder comprises 8 convolutional layers that progressively downsample the input cube to a bottleneck representation, while the decoder uses 8 transposed convolutional layers to upsample the bottleneck output to the target cube. The channel configuration (1+3) corresponds to the density field, $\delta_{\mathrm{m}}$, and the velocity components $(v_{x},v_{y},v_{z})$, enabling simultaneous reconstruction of these quantities as an optional feature. The computational steps are represented by distinct symbols, as illustrated in the lower right corner. All convolutional layers utilize $4\times 4\times 4$ filters with a stride of $2$.
@@ 3.2 UNet Architecture
We base our network architecture on the UNet style (Ronneberger et al., 2015), which is particularly well suited to reconstructing complex fields, due to its hierarchical convolutional layers and skip connections. These design features allow the model to capture multiscale spatial information effectively, which is critical for learning the mapping between the redshift-space galaxy density field and the target fields.
@@
  • Encoder Path: The encoder extracts features from the input field, $\delta_{\mathrm{g}}$, while progressively reducing the spatial resolution. Each block consists of a sequence: convolution, batch normalization, and LeakyReLU activation. Batch normalization is omitted in the first and last layers of the encoder, where the last layer uses ReLU activation instead of LeakyReLU.
  • Decoder Path: The decoder reconstructs the target field incrementally, by increasing spatial resolution. Each block consists of a sequence: transposed convolution, batch normalization, dropout (with a rate of 50%), and ReLU activation. Batch normalization is not applied in the first layer of the decoder, and the output layer employs a linear activation function.
  • Skip Connections: Skip connections link the encoder layers to their corresponding decoder layers, enabling the preservation of spatial details during downsampling and upsampling.
  • Output Layer: For the density field reconstruction, the output layer generates a single-channel output corresponding to $\delta_{\mathrm{m}}$. For the velocity field reconstruction, the output layer produces three channels, representing the velocity components $(v_{x},v_{y},v_{z})$.
In this architecture, the fields $\delta_{\mathrm{m}}$ and $\boldsymbol{v}_{\mathrm{m}}$ are reconstructed separately, due to memory limitations, while they could be reconstructed simultaneously, by designing a multichannel output for the network hardware permitting. A detailed comparison between the separately trained and unified models is provided in Appendix B.
Figure 6: Upper left: loss function for training $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ using 3000 epochs. The dark solid and blue dashed lines represent the loss functions from the training and validation samples, respectively. Upper right: comparison of the density distributions for the validation samples. The black solid line represents the result of the target validation samples. Different colored lines represent the predicted fields at epochs 200, 1000, 2000, and 3000, repectively. Lower left: grid-to-grid density relation between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$. The contour in each case represents threshold encompassing 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. The different colored curves correspond to different epochs, as indicated. Lower right: the evolution of the factor $f_{sc}$, defined as the product of the scatter variance and one minus the correlation scope between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as a function of epoch.
@@ 3.3 Training Process
As described in Section 2, we generate paired data samples $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ and $(\delta_{\mathrm{g}},\boldsymbol{v}_{\mathrm{m}})$ to train the networks. These samples represent the galaxy density field $\delta_{\mathrm{g}}$ paired with the true density field $\delta_{\mathrm{m}}$ or the velocity field $\boldsymbol{v}_{\mathrm{m}}$. The objective is to learn the mapping between the galaxy and dark matter fields.
The training process then minimizes the MSE loss function, which corresponds to the log-likelihood function in Equation (5). For practical implementation, the loss function is expressed as
@@
where $\phi^{\mathrm{pred}}_{\mathrm{m}}$ is the field being reconstructed and $\phi^{\mathrm{true}}_{\mathrm{m}}$ denotes the true target field. Specifically, $\phi_{\mathrm{m}}\equiv\delta_{\mathrm{m}}$ for density reconstruction and $\phi_{\mathrm{m}}\equiv\boldsymbol{v}_{\mathrm{m}}$ for velocity reconstruction.
To improve the training stability and optimize the network’s ability to capture variations across a wide range of values, we rescale the input fields before training. The overdensity field is rescaled using a logarithmic transformation:
@@
where the parameter $a=0.02$ ensures numerical stability, by preventing singularities for low-density regions ($\delta\approx-1$). Moreover, the $a$ value is optimized in our analysis to balance the reconstruction, enhancing the accuracy in high-density regions while preserving reliable performance in low-density areas. The velocity field is rescaled linearly as
@@
where the parameter $b=3000\,\mathrm{km}\,\mathrm{s}^{-1}$ normalizes the velocity values to a comparable scale, avoiding excessively large gradients during training. These rescaling transformations enhance the network’s convergence and improve its performance, by standardizing the input dynamic range.
@@
The network is trained from scratch using the Adam optimizer (Loshchilov & Hutter, 2017), which incorporates the adaptive learning rate. Each epoch processes five training samples, followed by two validation samples, used exclusively for model evaluation during training.
@@
When training neural networks, selecting an optimal stopping point is critical to balancing generalization and overfitting, particularly for fields with wide-ranging density distributions. As shown in the upper left panel of Figure 6, the validation loss converges around epoch 200, indicating an appropriate point for terminating training. However, relying solely on loss metrics may overlook the model performance across varying density regions, especially in high-density areas, which represent a small fraction of the total volume.
@@
To address this limitation, we compare the density distributions of the predicted and true fields (upper right panel of Figure 6). From epoch 200 to 2000, the disparities in high-density regions with $(1+\delta)>6$ progressively diminish, indicating significant improvements in recovering high-density structures even after loss convergence. However, as training progresses, predictions exhibit systematic underestimation at epoch 3000, revealing overfitting effects.
@@
However, the grid-to-grid relationship (lower left panel of Figure 6) supports the opposite trend. From epoch 200 to 2000, the scatter increases while maintaining minimal bias, indicating improved density distribution recovery at the cost of reduced correlation with the true field. This trade-off highlights the need for a quantitative criterion to identify the optimal training stage.
@@
We define the factor $f_{sc}$ as the product of the perpendicular variance, represented by the smaller eigenvalue of the covariance matrix of $(\vec{\delta}_{\mathrm{rec}},\vec{\delta}_{\mathrm{true}})$, and the absolute value of 1 minus the correlation slope, which is given by the slope of the eigenvector corresponding to the larger variance. As illustrated in the lower right panel of Figure 6, this factor reaches its minimum value at approximately epoch 1300 and increases significantly after epoch 2000. Consequently, halting the training process around epoch 1300 achieves a balance between high-density recovery and correlation retention, ensuring the reconstructed field accurately represents cosmological structures.
@@
These results underscore the necessity of carefully assessing the training stages to ensure the reliability of the model for cosmological inference. The methodology introduced here could be generalized to other datasets (e.g. Shi et al., 2024), offering a pathway to robust ML applications in astrophysical studies.
@@
Figure 7: Comparison of the projected density field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. The different panels correspond to the redshift-space galaxy distribution (input), the true dark matter density field $\delta_{\mathrm{true}}$ (target), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (output), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as labeled at the top of each panel.
@@
Figure 8: Left: grid-to-grid density relationship between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$, represented by the blue solid line. For comparison, the result for the galaxy density field is also plotted with the red dashed line. The three contours in each case represent thresholds encompassing 67%, 95%, and 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. Right: density distributions for galaxies (red dashed line), the reconstructed field (blue solid line), and the true field (black solid line). The error bars indicate the $\pm 1\sigma$ variance among the five test samples.
@@
Figure 9: The Same as Figure 8 but comparing the densities in the low-redshift (orange) and high-redshift (blue) region .
@@
Figure 10: Comparison of $P(k)$ (upper panel) and $P(k)$ (lower panel) ratios. The black solid line and blue dashed line represent the autopower spectrum of the true and reconstructed density fields respectively, while the green dotted-dashed line represents their cross-power spectra. For reference, the red dotted line shows the autopower spectra of the redshift-space galaxy field. The bottom panel shows the ratio between the reconstructed and true power spectra, with the error bars indicating the $\pm\sigma$ variance across the five test samples. The black solid error bars are computed using Eq. (9).
@@
Figure 11: The ratios $R_{\ell}(k)$ for the quadrupole (blue solid) and hexadecapole (orange dashed) power spectra.
@@ 4 Results
In this section, we present the reconstructed density and velocity fields derived from the mock DESI survey, by applying the resulting UNet model to five test samples. We also assess the reconstructed density field by validating its tidal field.
4.1 Reconstructing the density field
For visual comparison, Figure 7 shows the projected density field distribution in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. The Different panels correspond to the redshift-space galaxy distribution (upper left), the true dark matter density field $\delta_{\mathrm{true}}$ (upper right), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (lower left), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$ (lower right). Compared with the true field, the reconstructed density field reveals recognizable and consistent LSSs, including clusters, filaments, and voids, indicating a successful reconstruction overall. Additionally, the results demonstrate resilience to boundary effects, as structures near the survey edges are effectively recovered. However, the difference map highlights discrepancies in high-density regions, particularly around clusters, where the reconstructed field deviates from the target.
@@ 4.1 Reconstructing the density field
The right panel of Figure 8 shows the density distributions for galaxies (red dashed line), the reconstructed field (blue solid line), and the true field (black solid line). The error bars indicate the $\pm 1\sigma$ variance among the five test samples. After reconstruction, $\delta_{\mathrm{rec}}$ closely matches $\delta_{\mathrm{true}}$ in the regions where $1+\delta<7$, but a systematic underestimation occurs for $1+\delta>7$. This discrepancy likely arises from the imbalance in the number of grid cells between low- and high-density regions. As discussed in Section 3.3, the network initially prioritizes accurate predictions in low-density regions during early training stages, gradually improving high-density recovery at the cost of reduced correlation with the true field during later stages. The final model was selected by balancing scatter and correlation measures, which may contribute to the observed divergence. However, this issue affects only about 0.2% of the grids within the $1+\delta>7$ region. In comparison, the galaxy density distribution fails to trace the dark matter field accurately across both low- and high-density regimes.
@@
Additionally, Figure 9 presents a comparison between the low-redshift ($0.1\lesssim z\lesssim 0.2$) and high-redshift ($0.3\lesssim z\lesssim 0.4$) regions. In the left panel, the high-redshift region exhibits slightly larger scatter, highlighting the impact of lower galaxy density. In the right panel, the reconstructed density distribution underestimates high-density regions more significantly in the low-redshift region. This underestimation likely arises from higher galaxy clustering at low redshifts, which enhances nonlinear effects that the reconstruction method cannot fully correct, resulting in a suppression of high-density peaks. Overall, the method demonstrates robust performance, though further refinement is needed to better address nonlinear effects.
@@
To evaluate the clustering, we compare the power spectra of the reconstructed density fields with the true fields. In this study, random samples were not used to correct for the survey’s geometry or selection effects, as the focus is on evaluating the reconstruction’s fidelity through direct comparison. This highlights the quality of the reconstruction but does not represent the true underlying clustering.
@@
The top panel of Figure 10 illustrates the power spectrum comparisons. The black solid line and blue dashed line represent the autopower spectra of the true and reconstructed density fields, respectively. For reference, the red dotted line shows the autopower spectrum of the redshift-space galaxy density field. The bottom panel shows the ratio between the reconstructed and true power spectra, with the error bars indicating the $\pm\sigma$ variance across the five test samples. By considering the cosmic variance due to a limited number of Fourier modes for a finite-volume survey, we present the theoretical error (black error bar) for the power spectrum(e.g. Song et al., 2021) using
@@
Figure 12: Comparison between the true and reconstructed dark matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented.
To evaluate the concordance between the reconstructed and true density fields, we analyze the cross-correlation power spectrum. Unlike the autopower spectrum, which assesses the overall clustering amplitude, the cross-power spectrum directly measures how well the reconstructed field reproduces the phase and spatial distributions of the true field. High cross-correlation values indicate that the reconstruction effectively captures both large- and small-scale structures, making it a robust validation metric. In Figure 10, the green dashed-dotted line represents the cross-power spectrum of the true and reconstructed fields, as well as the ratios of the reconstructed cross-power spectrum to the true one. The results show that the ratios align with unity within the $1\sigma$ range on scales $k<0.1\,h\,\mathrm{Mpc}^{-1}$, indicating a strong agreement between the two fields. At smaller scales ($k>0.1\,h\,\mathrm{Mpc}^{-1}$), the ratios fall below unity. Specifically, the reductions in the cross-correlation power spectrum are approximately 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$. These reductions, though moderate, reflect challenges in accurately reconstructing high-density regions at smaller scales. Quantitatively, the average cross-correlation coefficient remains above 0.985 for $k<0.1\,h\,\mathrm{Mpc}^{-1}$, confirming the high fidelity of the reconstructed field for LSSs.
To evaluate the accuracy of the RSD corrections in the reconstructed density field, we compute the ratios of the multipole moments of the power spectra. These ratios compare the reconstructed and true power spectra, normalized by the difference between the redshift-space galaxy density power spectrum and the true power spectrum. The general formula for the ratio for any multipole $\ell$ is expressed as: